There is a number between 10 and 100: when multiplied by the digit on the left the product is 280: if the sum of the digits be multiplied by the same digit the product is 55: find the number.
Answer by sudhanshu_kmr(1152)


Let assume number is 10x+y i.e x at 10s place, y at unit place.
(10x+y)*x = 280
(x+y)*x = 55
55 = 5*11 ( only two factor) so, it is clear that x= 5
(x+y) = 11 so, y = 6

required number is : 56